# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ 2.4.3. Definition of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$
    The map $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}\rightarrow\mathsf{M}_{g,m}$ is representable by manifolds which are locally products of hyperplane complements in affine spaces; in particular, it is a submersion. The fibre product of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$ with a point over $\mathsf{M}_{g,m}$ is a product of $m$ products of hyperplane complements in complex vector spaces as defined in [DRT, Definition 1.3].
@@
We know from the above proposition that $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$ is a stack, which implies that the same is true for $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$.
Let us now prove the second point. Let $B\rightarrow\mathsf{M}_{g,m}$ be a manifold as in the statement, and let $B^{\prime}$ be a manifold. By definition (cf. section 4.1.6), the groupoid $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}\times_{\mathsf{M}_{g,m}}B(B^{\prime})$ has the following description.
  • Objects are triples $(\varphi\colon B^{\prime}\rightarrow B,(\pi^{\prime}\colon\Sigma^{\prime}\rightarrow B^{\prime},\underline{a}^{\prime},\underline{Q}^{\prime}),\tau)$, where $\tau\colon\Sigma\times_{B,\varphi}B^{\prime}\rightarrow\Sigma^{\prime}$ is an isomorphism commuting with the maps to $B^{\prime}$ and respecting marked points.
Therefore, we see that every object of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}\times_{\mathsf{M}_{g,m}}B(B^{\prime})$ is isomorphic to one of the form $(\varphi\colon B^{\prime}\rightarrow B,(\Sigma\times_{B}B^{\prime}\rightarrow B^{\prime},\varphi^{*}(\underline{a}),\underline{Q}^{\prime}),\mathrm{Id})$, where $\underline{Q}^{\prime}$ is a collection of irregular types on $\Sigma\times_{B}B^{\prime}$; furthermore, objects of this form have no automorphisms. In other words, the groupoid $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}\times_{\mathsf{M}_{g,m}}B(B^{\prime})$ is equivalent to $\mathscr{I\!\!T}^{\mathfrak{t},\leq p}_{B,a_{1},\ldots,a_{m}}(B)$.
Let us now deduce that $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$ is an analytic stack. Choose a surjective submersion $U\rightarrow\mathsf{M}_{g,m}$ from a complex manifold $U$. Then the fibre product $V=U\times_{\mathsf{M}_{g,m}}\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$ is a manifold, and the map $V\rightarrow\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$ is a surjective submersion, hence $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$ is an analytic stack. The same argument shows that the fact that $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$ is analytic follows from point $(4)$.
To prove $(4)$, take a map $B\rightarrow\mathsf{M}_{g,m}$ corresponding to an $m$-pointed genus $g$ Riemann surface $(\Sigma\rightarrow B,\underline{a})$. The fibre product $B\times_{\mathsf{M}_{g,m}}\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$ is representable by the product (over $B$) of the spaces representing $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}_{i}}_{B,a_{i}}$, for $1\leq i\leq m$. We have seen in section 2.3.5 that each $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\leq\mathbf{d}_{i}}_{B,a_{i}}$ is representable by a manifold over $B$ which, locally on $B$, is an intersection of the manifolds
(2.4.6.1) $B\times\left(\prod_{1\leq j\leq d_{\alpha,i}}\mathfrak{t}\times\prod_{d_{\alpha,i}<j\leq p}\mathrm{ker}(\alpha)\right)\,,\qquad\alpha\in\Phi\,.$
Therefore, the functor $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}_{i}}_{B,a_{i}}\colon\mathrm{Man}_{\mathbf{C}}\rightarrow\mathrm{Sets}$ of irregular types at $a_{i}$ with root order $\mathbf{d}_{i}$ is representable by a manifold which, locally on $B$, is the open submanifold of the intersection of the manifolds in (2.4.6.1) whose $j$-th component, for $1\leq j\leq p$, is the product of $B$ and
$\bigcap_{d_{\alpha,i}<j}\mathrm{ker}(\alpha)\cap\bigcap_{d_{\alpha,i}=j}(\mathfrak{t}\smallsetminus\mathrm{ker}(\alpha))\,.$
The above manifold coincides with the one defined in [DRT, Definition 1.3], hence the proof of $(4)$ is complete. ∎
2.4.7 Corollary.
If $\mathsf{M}_{g,m}$ is (representable by) a manifold then $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$ is a manifold as well, and the map $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}\rightarrow\mathsf{M}_{g,m}$ is a holomorphic submersion.
3. Wild mapping class groups
In this section we define and study the fundamental groups of the stacks of admissible (families of) wild Riemann surfaces, called wild mapping class groups.
3.1. Definition and basic properties
Thanks to Theorem 2.4.6 we know that $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$ is an analytic stack. To such an object, once a base point is chosen, one can attach a well-defined (topological) fundamental group, cf. Definition 4.2.5.

### v2

@@ 4.2.3. The case $m=1$
    for $p=1$, an element $(s,r)$ acts by sending $A_{1}$ to $rA_{1}$. Indeed, $rA_{1}z+sA_{1}$ differs from $rA_{1}z$ by a term holomorphic at infinity;
  2. (2)
    for $p\geq 2$ instead we look at the subleading coefficient:
    $(s,r)\cdot Q-r^{p}A_{p}z^{p}=(pr^{p-1}sA_{p}+r^{p-1}A_{p-1})z^{p-1}+\sum_{j=1}^{p-2}A^{\prime}_{j}z^{j}.$
In the first case $\mathbf{C}\subset G_{1}$ acts trivially on $\mathscr{I\!\!T}^{\mathfrak{t},\leq 1}$; in particular, stabilisers in $G_{1}$ of points of $\mathscr{I\!\!T}^{\mathfrak{t},\leq 1}$ are never finite. Thus, given $\mathbf{d}=(d_{\alpha})_{\alpha\in\Phi}$ with $0\leq d_{\alpha}\leq 1$, for any $p\geq 1$ the space $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}$, if non-empty, contains points with stabilisers of infinite order.
In the second case the identity $(s,r)\cdot Q=Q$ forces $r$ to be a $p$-th root of unity, if $A_{p}\neq 0$. Furthermore, for each possible $r$, there is at most one $s$ such that $(pr^{p-1}sA_{p}+r^{p-1}A_{p-1})=A_{p-1}$. Hence, stabilisers of points in strata $\mathsf{WM}_{0,1}^{\mathfrak{t},\leq p,\mathbf{d}}$ are finite whenever $\max(\mathbf{d})\geq 2$.
More precisely, in case (2) the stack $\mathsf{WM}_{0,1}^{\mathfrak{t},\leq p,\mathbf{d}}$ is isomorphic to a locally closed substack of a weighted projective stack. To see this, choose $\alpha\in\Phi$ such that $d_{\alpha}\geq 2$, and let $Z_{\alpha}\subset\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}$ be the submanifold of irregular types such that $\alpha(A_{d_{\alpha}-1})=0$. Consider the map $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}\rightarrow Z_{\alpha}$ sending $Q$ to $(s(Q),1)\cdot Q$, where $s(Q)=-\frac{\alpha(A_{d_{\alpha}-1})}{p\alpha(A_{d_{\alpha}})}$. It induces a map $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}/\mathbf{C}\rightarrow Z_{\alpha}$ inverse to the one induced from the inclusion $Z_{\alpha}\subset\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}$. Therefore the map $Z_{\alpha}\rightarrow\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}/\mathbf{C}$ is a biholomorphism, hence we find that $\mathsf{WM}_{0,1}^{\mathfrak{t},\leq p,\mathbf{d}}=[Z_{\alpha}/\mathbf{C}^{\times}]$. The group $\mathbf{C}^{\times}$ acts on $Z_{\alpha}$ as follows: an element $r\in\mathbf{C}^{\times}$ sends an irregular type with coefficients $(A_{1},\ldots,A_{p})$ to the irregular type with coefficients $(rA_{1},\ldots,r^{p}A_{p})$. Therefore $Z_{\alpha}/\mathbf{C}^{\times}]$ is locally closed substack of a weighted projective stack.
4.2.4. The case $m=2$
Fix $\underline{p}=(p_{1},p_{2})\in\mathbf{Z}_{\geq 0}^{2}\smallsetminus\{(0,0)\}$, and tuples $\underline{\mathbf{d}}=(\mathbf{d}_{1},\mathbf{d}_{2})$ with $\mathbf{d}_{i}=(d_{\alpha,i})_{\alpha\in\Phi}$ such that $0\leq d_{\alpha,i}\leq p_{i}$. Let us assume that the integers $d_{\alpha,i}$ are not all equal to 0. Let $\mathscr{I\!\!T}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}=\mathscr{I\!\!T}^{\mathfrak{t},\leq p_{1},\mathbf{d}_{1}}\times\mathscr{I\!\!T}^{\mathfrak{t},\leq p_{2},\mathbf{d}_{2}}$. We identify this product with a moduli space of irregular types on $(\mathbf{P}^{1},(0,\infty))$, and endow it with the resulting action of $G_{2}=\mathbf{C}^{\times}$. Arguing as in the previous section one shows that
$\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}=[\mathscr{I\!\!T}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}/G_{2}].$
Therefore, we obtain an exact sequence
$\mathbf{Z}\simeq\pi_{1}(G_{2})\rightarrow\pi_{1}(\mathscr{I\!\!T}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}})\rightarrow\pi_{1}(\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}})\rightarrow 1,$
which shows in particular that the middle term is independent of the choice of $\underline{p}$ (subject to the condition $p_{i}\geq\max(\mathbf{d}_{i}))$. Note that one could also have obtained an exact sequence as above from the fibration $\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}\rightarrow\mathsf{M}_{0,2}$, using that $\pi_{1}(\mathsf{M}_{0,2})$ is trivial and $\pi_{2}(\mathsf{M}_{0,2})=\mathbf{Z}$.
The action of $G_{2}$ on $\mathscr{I\!\!T}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ always has finite stabilisers. Indeed, a point of $\mathscr{I\!\!T}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ is of the form $(Q_{1},Q_{2})\neq(0,0)$, where $Q_{1}=\sum_{j=1}^{p_{1}}A_{j,1}z^{-j}$ and $Q_{2}=\sum_{j=1}^{p_{2}}A_{j,2}z^{j}$. The action of $r\in\mathbf{C}^{\times}$ is obtained replacing $z$ by $rz$, hence sending $A_{j,1}$ to $r^{-j}A_{j,1}$ and $A_{j,2}$ to $r^{j}A_{j,2}$; therefore, stabilisers are finite groups of roots of unity.
If $\underline{p}$ is of the form $\underline{p}=(0,p_{2})$ with $p_{2}>0$ then $\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ is the quotient of $\mathscr{I\!\!T}^{\mathfrak{t},\leq p_{2},\mathbf{d}_{2}}$ by the action sending $A_{j,2}$ to $r^{j}A_{j,2}$. As $\mathbf{d}_{2}$ is not identically 0 (because we are assuming that the $d_{\alpha,i}$ are not all 0) one of the $A_{j,2}$ is necessarily non-zero. Hence, the stack $\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ is a locally closed substack of a weighted projective stack. For general $\underline{p}$, the stack $\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ is representable over a weighted projective stack. Indeed, we may assume without loss of generality that $\mathbf{d}_{2}$ is not identically 0; let $\underline{p}^{\prime}=(0,p_{2})$ and $\underline{\mathbf{d}}^{\prime}=(0,\mathbf{d}_{2})$. Forgetting the first irregular type we obtain a map $\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}\rightarrow\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p}^{\prime},\underline{\mathbf{d}}^{\prime}}$ which is representable: indeed, the same argument as in the proof of Theorem 2.4.6 shows that the fibre product of $\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ with a manifold $B\rightarrow\mathsf{WM}_{0,2}^{\mathfrak{t},\leq\underline{p}^{\prime},\underline{\mathbf{d}}^{\prime}}$ is isomorphic to $\mathscr{I\!\!T}^{\mathfrak{t},\leq p_{1},\mathbf{d}_{1}}_{B,a_{1}}$.
4.2.5. Some isomorphisms between moduli spaces of wild Riemann surfaces of genus 0
Fix integers $n\geq 0$ and $m\geq 0$, and let $\mathfrak{t}_{n+1}\subset\mathfrak{sl}_{n+1}$ be the subalgebra of diagonal matrices of trace zero. Let $\mathfrak{t}_{n+1}^{\mathrm{reg}}\subset\mathfrak{t}_{n+1}$ be the subspace of matrices with distinct entries. We will construct an isomorphism
(4.2.5.1) $\mathsf{WM}_{0,m+2}^{\mathfrak{t}_{n+1},\leq\underline{p},\underline{\mathbf{d}}}\xrightarrow{\sim}\mathsf{WM}_{0,n+2}^{\mathfrak{t}_{m+1},\leq\underline{p^{\prime}},\underline{\mathbf{d}^{\prime}}},$
where $\underline{p}$ (resp. $\underline{p^{\prime}}$) consists of one 1 followed by $m+1$ (resp. $n+1$) 0, and $\mathbf{d}_{1}$, $\mathbf{d}_{1}^{\prime}$ are identically equal to 1. This rests on the following elementary lemma, where $X_{n}$ denotes the complement of the diagonals in $(\mathbf{C}^{\times})^{n}$.
4.2.6 Lemma.
For $n\geq 1$, if $\mathrm{diag}(b_{1},\ldots,b_{n+1})$ belongs to $\mathfrak{t}_{n+1}^{\mathrm{reg}}$ then $(b_{2}-b_{1},\ldots,b_{n+1}-b_{1})$ belongs to $X_{n}$. The resulting map $\varphi_{n}:\mathfrak{t}_{n+1}^{\mathrm{reg}}\rightarrow X_{n}$ is a biholomorphism.
@@ 4.2.5. Some isomorphisms between moduli spaces of wild Riemann surfaces of genus 0
The map sending $(x_{1},\ldots,x_{n})$ to $(b_{1},\ldots,b_{n+1})$ where $b_{1}=-\frac{1}{n+1}(x_{1}+\ldots+x_{n})$ and $b_{i}=b_{1}+x_{i-1}$ for $i\geq 2$ is the inverse of $\varphi_{n}$. ∎
We obtain a biholomorphism
(4.2.6.1) $(\varphi_{n},\varphi_{m}^{-1}):\mathfrak{t}_{n+1}^{\mathrm{reg}}\times X_{m}\rightarrow X_{n}\times\mathfrak{t}_{m+1}^{\mathrm{reg}}.$
The left-hand side is the moduli space of $(m+2)$-tuples of distinct points on $\mathbf{P}^{1}$ of the form $(a_{1}=0,a_{2},\ldots,a_{m+1},a_{m+2}=\infty)$, together with a $\mathfrak{t}_{n+1}$-valued irregular type at 0 of pole order one and root order one at each root. The right-hand side has a similar modular interpretation, with $m$ and $n$ switched. The group $\mathbf{C}^{\times}$ of automorphisms of $\mathbf{P}^{1}$ fixing 0 and $\infty$ acts on the two spaces in (4.2.6.1): the action of $\alpha\in\mathbf{C}^{\times}$ on $X_{m}$ pulling back points sends $(x_{1},\ldots,x_{m})\in X_{m}$ to $(\alpha^{-1}x_{1},\ldots,\alpha^{-1}x_{m})$, and the action pulling back irregular types sends $\frac{Q}{z}$ to $\frac{Q}{\alpha z}$, hence it is given by multiplication by $\alpha^{-1}$ on $\mathfrak{t}_{n+1}^{\mathrm{reg}}$. For the same reason $\alpha$ acts by multiplication by $\alpha^{-1}$ on the right-hand side of (4.2.6.1), hence the map in (4.2.6.1) induces the isomorphism in (4.2.5.1).
4.3. Genus one examples
We have $\mathsf{M}_{1,1}=[\mathbf{H}/\mathrm{SL}_{2}(\mathbf{Z})]$, where $\mathbf{H}=\{\tau\in\mathbf{C}\mid\mathrm{Im}(\tau)>0\}$ and $\mathrm{SL}_{2}(\mathbf{Z})$ acts on it via Möbius transformations.
4.3.1.
Over $\mathbf{H}$ we have an elliptic curve (i.e., a family of compact 1-pointed Riemann surfaces of genus one)
$\mathcal{E}=(\mathbf{H}\times\mathbf{C})/\mathbf{Z}^{2},\;\sigma\colon\mathbf{H}\rightarrow\mathcal{E},$
where $(r,s)\in\mathbf{Z}^{2}$ acts sending $(\tau,z)$ to $(\tau,z+r\tau+s)$, and $\sigma(\tau)=(\tau,0)$. Therefore, the fibre $\mathcal{E}_{\tau}$ of $\mathcal{E}$ over $\tau\in\mathbf{H}$ is the torus $\mathbf{C}/(\mathbf{Z}+\tau\mathbf{Z})$ with the marked point 0. The action of $\mathrm{SL}_{2}(\mathbf{Z})$ on $\mathbf{H}$ lifts to an action on $\mathcal{E}$, induced by
(4.3.1.1) $\mathrm{SL}_{2}(\mathbf{Z})\times(\mathbf{H}\times\mathbf{C})$ $\rightarrow\mathbf{H}\times\mathbf{C}$
$\left(\begin{pmatrix}a&b\\ c&d\end{pmatrix},(\tau,z)\right)$ $\mapsto\left(\frac{a\tau+b}{c\tau+d},\frac{z}{c\tau+d}\right).$
The coordinate $z$ is a local coordinate around each $\sigma(\tau)$, hence irregular types with a pole at $\sigma$ of order at most $p$ can be written as
(4.3.1.2) $\sum_{j=1}^{p}z^{-j}A_{j}(\tau),$
where each $A_{i}\colon\mathbf{H}\rightarrow\mathfrak{t}$ is a holomorphic function. Therefore, we see that $\widetilde{WT}_{1,1}^{\mathfrak{t},\leq p}=\mathsf{WM}_{1,1}^{\mathfrak{t},\leq p}\times_{\mathsf{M}_{1,1}}\mathbf{H}$ is isomorphic to $\mathbf{H}\times\mathfrak{t}^{p}$. Furthermore, it comes with an action of $\mathrm{SL}_{2}(\mathbf{Z})$ which can be written down explicitly using (4.3.1.1), and determines the vector bundle $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq p}\rightarrow\mathsf{M}_{1,1}$. For instance, for $p=1$ the $\mathrm{SL}_{2}(\mathbf{Z})$-action on irregular types is given by
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\cdot\frac{A_{1}(\tau)}{z}=(c\tau+d)\frac{A_{1}(\frac{a\tau+b}{c\tau+d})}{z}.$
Therefore, the stack $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1}$ is the quotient of $\widetilde{WT}_{1,1}^{\mathfrak{t},\leq 1}\simeq\mathbf{H}\times\mathfrak{t}$ by the $\mathrm{SL}_{2}(\mathbf{Z})$-action given by
$\left(\begin{pmatrix}a&b\\ c&d\end{pmatrix},(\tau,A)\right)\mapsto\left(\frac{a\tau+b}{c\tau+d},\frac{1}{c\tau+d}A\right).$
In other words, $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1}$ is isomorphic to the product (over $\mathsf{M}_{1,1}$) of $r$ copies of the line bundle $\mathcal{L}^{-1}=(\mathbf{H}\times\mathbf{C})/\mathrm{SL}_{2}(\mathbf{Z})$, where the action of $\mathrm{SL}_{2}(\mathbf{Z})$ is given by (4.3.1.1). Its inverse $\mathcal{L}$ is isomorphic to the Hodge bundle on $\mathsf{M}_{1,1}$ (cf. [Hai11, § 5.4]), which generates the Picard group of $\mathsf{M}_{1,1}$, isomorphic to $\mathbf{Z}/12\mathbf{Z}$ [Hai11, Theorem 6.6] (see also [Mum65]).
In particular, $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1}\rightarrow\mathsf{M}_{1,1}$ is a non-trivial vector bundle.
4.3.2. The case $p=1,\mathfrak{g}=\mathfrak{sl}_{2}$
Let $p=1$ as above. Let us take $\mathfrak{t}$ to be the diagonal matrices in $\mathfrak{g}=\mathfrak{sl}_{2}$, and let $\alpha$ be the usual choice of a positive root. Let $B\rightarrow\mathsf{M}_{1,1}$ be a map corresponding to an elliptic curve $(\pi\colon\Sigma\rightarrow B,\sigma)$. By Theorem 2.4.6 the fibre product $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1}\times_{\mathsf{M}_{1,1}}B$ is isomorphic to the vector bundle attached to the locally free sheaf $\pi_{*}(\mathcal{I}_{\sigma}^{-1}/\mathcal{O}_{\Sigma})\simeq\sigma^{*}(\mathcal{I}_{\sigma}^{-1}/\mathcal{O}_{\Sigma})$, which is the inverse of $\sigma^{*}(\mathcal{I}_{\sigma}/\mathcal{I}_{\sigma}^{2})\simeq\sigma^{*}(\Omega^{1}_{\Sigma/B})$ (where the isomorphism follows from the conormal exact sequence). We see in particular once again that $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1}$ corresponds to the inverse of the Hodge bundle over $\mathsf{M}_{1,1}$.
Let us now consider the substack $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1,1}\subset\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1}$ parametrising admissible families of wild Riemann surfaces with pole order 1 at $\alpha$. In other words, this is the complement of the zero section inside $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1}$. Explicitly, we have $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1,1}\simeq(\mathbf{H}\times\mathbf{C}^{\times})/\mathrm{SL}_{2}(\mathbf{Z})$, where the action is given by the formula in (4.3.1.1). Note that this action is free, and the stack $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1,1}$ is a complex manifold. The short exact sequence in section 3.1.3 reads
$1\rightarrow\mathbf{Z}\rightarrow\mathsf{W}\Gamma_{1,1}^{\mathfrak{t},1}\rightarrow\mathrm{SL}_{2}(\mathbf{Z})\rightarrow 1.$
Finally, the map $\mathbf{H}\times\mathbf{C}^{\times}\rightarrow\mathbf{H}\times\mathbf{C}^{\times}$ sending $(\tau,z)$ to $(\tau,z^{-1})$ induces an isomorphism between $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq 1,1}$ and the complement $\mathcal{L}^{\times}$ of the zero section inside the Hodge bundle. The fundamental group of $\mathcal{L}^{\times}$ is isomorphic to the braid group $B_{3}$ [Hai11, Corollary 8.3], therefore we find that
$\mathsf{W}\Gamma_{1,1}^{\mathfrak{t},1}\simeq B_{3}.$
5. Stacks of wild algebraic curves
In this section we explain how to adapt the above discussion to define algebraic stacks $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ of (admissible families of) wild algebraic curves; this is straightforward, hence we will be brief. Our main goal is to determine which of the stacks $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ are Deligne–Mumford, confirming the expectation of [BDR25] in the untwisted case. We will work in the category $\mathrm{Sch}_{\mathbf{C}}$ of schemes over $\mathrm{Spec}(\mathbf{C})$, and $S$ will always denote such a scheme. If $\mathfrak{g},\mathfrak{t}$ and the roots $\alpha$ are defined over a field $F$, the construction of $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ works over $\mathrm{Spec}(F)$.
5.1. Wild algebraic curves
By a curve over $S$ of genus $g$ we mean a smooth, proper morphism of schemes $\pi\colon C\rightarrow S$ of relative dimension one, such that all the fibres are geometrically connected of genus $g$. Note that $C/S$ is a smooth curve in the sense of [KM85, §(1.2.1)]. By [KM85, Lemma 1.1.2], if $\sigma\colon S\rightarrow C$ is a section then the image $\sigma(S)$ is an effective Cartier divisor, i.e. its ideal sheaf $\mathcal{I}_{\sigma}$ is invertible, generated locally by a nonzerodivisor. Therefore, we can define $\mathscr{T}_{C,\sigma}^{\leq k}$, for $k\geq 0$, as in section 2.2. The results in the rest of section 2.2, which rely on the purely algebraic Lemma 2.2.5, carry over to this setting.
We define an irregular type (on $S$) with pole at $\sigma$ of order at most $k$ to be a global section of the sheaf $\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{C,\sigma}^{\leq k}$. The functor $\mathscr{I\!\!T}^{\mathfrak{t},\leq k}_{S,\sigma}$ from $S$-schemes to sets sending $S^{\prime}\rightarrow S$ to $\Gamma(C^{\prime},\mathfrak{t}\otimes\mathscr{T}_{C^{\prime},\sigma^{\prime}}^{\leq k})$, where $C^{\prime}=C\times_{S}S^{\prime}$ and $\sigma^{\prime}$ is the pullback of $\sigma$, is representable by a vector bundle over $S$, denoted by the same symbol. This is shown as Proposition 2.3.4 using [Sta18, Tag 02KG] to prove the analogue of Lemma 2.2.8, and [GW20, Proposition 11.3, §11.4].